Salaries of 32 college graduates who took a statistics course in college have a mean, x, of $64,600. Assuming a standard deviation, o, of $14,236, construct a 90% confidence interval for estimating the population mean . Click here to view a t distribution table Click here to view page 1 of the standard normal distribution table. Click here to view pag hal «D < μ<s[ ] (Round to the nearest integer as needed.)
Answer #1

SOLUTION given: 1-64600, σ-1423 6, n-32, c 0.90 formula: X tan Za * for 90% confidence level, Za = 1.645 (from z-table) 64600 1.645 * = 64600 ± 4140 Lower bound 64600-4140 60460 Upper bound = 64600 + 4140 = 68740 $60460 < μ < $68740
Answer #2

Salaries of college graduates who took a statistics course in college have a mean, , of  Assuming a standard deviation, , of $1, construct a % confidence interval for estimating the population mean .


source: person
answered by: kadie